# A Fractional Calculus Framework for Open Quantum Dynamics: From Liouville to Lindblad to Memory Kernels

### v1

@@ Abstract
Open quantum systems exhibit dynamics ranging from purely unitary evolution to irreversible dissipative relaxation. The Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation uniquely characterizes Markovian dynamics that are completely positive and trace-preserving (CPTP), yet many physical systems display non-Markovian features such as algebraic relaxation and coherence backflow beyond the reach of semigroup evolution. Fractional calculus provides a natural framework for describing such long-memory behavior through power-law temporal kernels introduced by fractional time derivatives. Here we establish a unified hierarchy that embeds fractional quantum master equations within the broader landscape of open system dynamics. The fractional master equation forms a structured subclass of memory-kernel models, reducing to the GKSL form at unit fractional order. Through Bochner–Phillips subordination, fractional evolution is expressed as an average over Lindblad semigroups weighted by a power-law waiting-time distribution. This construction ensures physical consistency, explains the algebraic origin of long-time decay, and bridges unitary, Markovian, and structured non-Markovian regimes. The resulting framework positions fractional calculus as a rigorous and unifying language for quantum dynamics with intrinsic memory, enabling new directions for theoretical analysis and quantum simulation.
@@ I Introduction
In this work, we develop a unified theoretical framework that places fractional dynamics within the broader hierarchy of quantum open system formalisms. We show that fractional master equations form a structured subclass of non-Markovian QMEs, coinciding with standard Lindblad dynamics in the limit of unit fractional order. This perspective reveals that fractional differentiation introduces memory at the level of the generator rather than through explicit environmental coupling, bridging the conceptual gap between purely unitary and dissipative descriptions.
@@ II.1 Preliminaries
Liouville $\alpha=1,\ \lambda=0$ Unitary, reversible Markovian Lindblad $\alpha=1,\ \lambda>0$ CPTP semigroup Non-Markovian QME $\alpha\!\neq\!1,\ \lambda>0$ Not always CPTP Fractional Lindblad $\prescript{C}{}{D}_{t}^{\alpha}\rho=\lambda\,\mathcal{L}\rho$ $0<\alpha<1,\ \lambda>0$ CPTP, power-law memory $\lambda\downarrow,~\alpha\uparrow$$\alpha\uparrowgeneral$ memory
Figure 1: Schematic hierarchy of quantum dynamical regimes unified by fractional subordination. The parameters $\alpha$ (memory order) and $\lambda$ (dissipation strength) interpolate smoothly between unitary Liouville dynamics ($\alpha{=}1,\lambda{=}0$), Markovian Lindblad semigroups ($\alpha{=}1,\lambda{>}0$), and general memory–kernel quantum master equations ($\alpha{\neq}1,\lambda{>}0$). Fractional or subordinated Lindblad dynamics ($0{<}\alpha{<}1,\lambda{>}0$) form the CPTP bridge connecting these regimes.
@@
III Fractional Master Equation
@@
III.1 Equivalence to convolution master equations
@@
If $\mathcal{L}\rho(0)=0$ (for instance, when $\rho(0)$ is a fixed point of $\mathcal{L}$), the boundary term vanishes and Eq. (11) reduces to a pure convolution master equation Chruscinski2017; Tarasov2021.
@@
III.2 Mittag–Leffler solution and physical interpretation
@@ III.3 Fractional Subordination
The kernel satisfies $f_{\alpha}(u,t)\geq 0$ and $\int_{0}^{\infty}f_{\alpha}(u,t)~du=1$, which makes the evolution (17) represents a convex mixture of CPTP semigroups $e^{u\mathcal{L}}$ (detailed in Appendix C). Specifically, the subordination integral in Eq. (17) represents a probabilistic construction of fractional dynamics: the inverse–stable density $f_{\alpha}(u,t)\!\geq 0$ acts as a genuine probability distribution for the random operational time $u$ of an underlying stochastic clock. Consequently, $\Phi_{\alpha}(t)$ can be interpreted as the statistical mixture $\Phi_{\alpha}(t)=\mathbb{E}_{U(t)}[e^{U(t)\mathcal{L}}]$, i.e., an ensemble average of standard Markovian (Lindblad) evolutions sampled at random times $U(t)$ with power–law waiting statistics. This probabilistic representation ensures that complete positivity is preserved and explains how algebraic relaxation emerges naturally from the statistics of random operational time.
The representation (17) also has an important algorithmic interpretation. Because $e^{u\mathcal{L}}$ admits standard quantum–trajectory and quantum–jump unravelings, the subordination formula shows that any Markovian trajectory method can be extended to fractional, non-Markovian dynamics simply by replacing the physical time by a random clock $U(t)$ drawn from $f_{\alpha}(u,t)$. Thus, fractional dynamics corresponds to sampling ordinary Lindbladian trajectories at random operational times, with the memory effects arising entirely from the statistics of $U(t)$.
@@
IV Numerical Demonstration and Discussion
We work in the exactly solvable pure–dephasing limit of the spin–boson model,
@@ IV Numerical Demonstration and Discussion
with independent boson modes $\{b_{k}\}$ in thermal equilibrium at inverse temperature $\beta$. Since $H,\sigma_{z}]=0$, populations are constants of motion; only the off–diagonal element (coherence) decays. If we assume an initially fractorized state for the total spin-boson system
@@
as used in standard treatments of dissipative two-level systems and spin–boson dynamics Grifoni1998; Weiss2012. Table 1 provides the asymptotic scaling for the coherence that smoothly interpolates between Gaussian short-time decay, stretched-exponential-like intermediate behavior, and algebraic or saturating long-time tails, depending on $s$ and $t$ Leggett1987; Grifoni1998; Weiss2012; deVega2017. The real-time propagation of the dephasing functional $Q(t)$ and corresponding coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for $t\leq 0.2/\omega_{c}$ and $t\geq 5/\omega_{c}$ and three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are shown in Figure 2. In the weak-coupling and pure-dephasing limit, the reduced coherence $u(t)$ obeys a Nakajima–Zwanzig (NZ) convolution equation Nakajima1958; Zwanzig1960,
@@
coincides with the Gaussian-bath result in Eq. (25), confirming that $Q(t)$ fully characterizes the non-Markovian dephasing kernel.
@@
$Q(t)$ $\mathcal{O}\left(\frac{1}{2}\eta\,\Gamma(\chi+1)\,\omega_{c}^{2}\,t^{2}\right)$ $\mathcal{O}\left(C_{\chi}(\omega_{c})\,t^{1-\chi}\right)$ $\mathcal{O}\left(\frac{\eta}{2}\ln(\omega_{c}^{2}t^{2})\right)$ $\mathcal{O}\left(Q_{\infty}-D_{\chi}\,t^{-(\chi-1)}\right)$
@@
IV.1 Why Markovian dephasing misses $e^{-Q(t)}$?

### v2

@@ Abstract
Open quantum systems exhibit dynamics ranging from unitary evolution to irreversible dissipation. While the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation uniquely characterizes Markovian CPTP evolution, many physical platforms display non-Markovian features such as algebraic relaxation and coherence backflow. Fractional calculus provides a natural way to model such long-memory behavior through power-law temporal kernels introduced by fractional time derivatives. Here we develop a unified framework that embeds fractional master equations within the broader hierarchy of open-system formalisms. The fractional equation forms a structured subclass of memory-kernel models, reduces to the Lindblad form at unit order, and, through Bochner–Phillips subordination, admits a CPTP representation as an average over Lindblad semigroups. Its resolvent structure further connects fractional dynamics to established non-Markovian approaches, including Nakajima–Zwanzig kernels and hierarchical equations of motion, providing a compact surrogate for long-memory effects. This formulation positions fractional calculus as a rigorous and practical language for quantum dynamics with intrinsic memory, supporting both analytical insight and efficient quantum simulation.
@@ I Introduction
Addressing these questions requires placing fractional dynamics within the broader QME landscape. Although numerically exact open-system solvers have advanced rapidly, QMEs themselves remain central because they provide analytical insight, scalable reduced descriptions, and generator-level models that interface naturally with quantum simulation and control frameworks. Clarifying where fractional dynamics sit within this landscape therefore involves relating them to existing QME-based approaches and understanding how they complement, rather than compete with, exact computational methods.
A wide variety of non-Markovian generalizations of the GKSL equation have been proposed. Projection-operator methods such as the Nakajima–Zwanzig (NZ) equation Nakajima1958; Zwanzig1960 introduce memory kernels through explicit system–bath couplings, while the time-convolutionless (TCL) formulation yields time-dependent generators that may lose CP-divisibility Breuer2002; Breuer2016. Generalized Lindblad approaches (e.g., Ref. Breuer2007GeneralizedLindblad) incorporate environmental correlations via correlated projection operators or extended state spaces, and typically retain a first-order time derivative but with time-dependent decay rates. Microscopic approaches such as hierarchical equations of motion (HEOM) HEOM_1; HEOM_2; HEOM_3; HEOM_4 provide numerically exact descriptions for Gaussian environments but at a high computational cost. These formalisms collectively establish a rich landscape of non-Markovian models, and recent years have seen substantial progress in improving their numerical efficiency, for example through tensor-network representations Chen:2025to; Yan:2021tg. Nevertheless, even in these advanced implementations, they generally lack the simplicity of a closed-form CPTP propagator or an analytically transparent interpolation between Markovian and memory-dominated regimes.
Beyond projection-operator and time-local master equations, many non-Markovian methods arise from influence-functional formalisms, including the quasi-adiabatic propagator path integral (QUAPI) Makri1995QUAPI1; Makri1995QUAPI2, multiconfiguration time-dependent Hartree (MCTDH) approaches to open-system dynamics Thoss2001MCTDH; Wang2003MCTDHReview, and related path-integral techniques. These frameworks provide controlled microscopic descriptions of dissipative dynamics but typically involve large tensor-network objects or exponential memory growth as temporal correlations accumulate. In contrast, the fractional approach developed here provides a compact, closed-form surrogate for environments with slow or algebraically decaying correlations. While it does not reproduce the full influence-functional structure, it captures the emergent memory behavior in a generator-level deformation of the Lindblad semigroup. Thus, fractional dynamics offer an analytically transparent complement to computationally intensive influence-functional methods, providing a reduced model that summarizes the dominant memory features encoded in more detailed path-integral treatments.
The above considerations motivate the development of a framework that situates fractional evolution within the established hierarchy of open-system formalisms while retaining the interpretability of generator-based descriptions. In this work, we develop a unified theoretical framework that places fractional dynamics within the broader hierarchy of quantum open system formalisms. We show that fractional master equations form a structured subclass of non-Markovian QMEs, coinciding with standard Lindblad dynamics in the limit of unit fractional order. This perspective reveals that fractional differentiation introduces memory at the level of the generator rather than through explicit environmental coupling, bridging the conceptual gap between purely unitary and dissipative descriptions.
Our key observation is that the fractional master equation admits a resolvent-level representation $(s^{\alpha}\mathbb{I}-\mathcal{L})^{-1}$ that can be interpreted through Bochner–Phillips subordination Bochner1955; Feller1971 as a convex mixture of Lindblad semigroups evaluated at Lévy-distributed operational times. This structure guarantees complete positivity for any GKSL generator and provides a closed analytical solution in terms of the Mittag–Leffler function. Moreover, we show that the fractional resolvent is algebraically connected to the Laplace-resolvent forms appearing in the NZ equation (long-tailed kernels), HEOM (self-energy corrections), and time-local generalized Lindblad approaches (time-dependent rates). In this sense, fractional dynamics emerge not as a phenomenological alteration of quantum mechanics but as a mathematically precise deformation of the Lindblad semigroup that captures algebraic memory in a compact, two-parameter form.
@@ II.1 Preliminaries
Remark 1 (Physical meaning of the fractional order $\alpha$.).
The fractional order $0<\alpha<1$ has a clear operational interpretation within the renewal–process framework underlying Eq. (6). In Bochner–Phillips subordination, the physical time $t$ is replaced by a random operational time $U(t)$ associated with an inverse–stable subordinator Bochner1955; Feller1971; Schilling2012Bernstein; BaeumerMeerschaert2001. The corresponding renewal process has a heavy–tailed waiting–time density $\psi_{\alpha}(\tau)\sim\tau^{-1-\alpha}$, a structural consequence of the Lévy measure of the underlying $\alpha$-stable subordinator METZLER20001; MainardiScalas2004; MeerschaertSikorskii2012. The parameter $\alpha$ therefore controls the heaviness of the waiting-time tail and thus the strength and persistence of memory in the stochastic time change. When $\alpha=1$, the waiting times become exponential and one recovers the Markovian Lindblad semigroup; when $\alpha<1$, the broad, power-law waiting statistics generate non-Markovian dynamics with Mittag–Leffler relaxation, stretched-exponential transients, and long-time algebraic tails.
In this sense, $\alpha$ is not a phenomenological constant but a structural memory exponent: it determines how far the dynamics deviate from Markovian behavior and how quickly the influence of past states decays. In microscopic settings such as the pure-dephasing spin–boson model, the value of $\alpha$ can be related to properties of the bath correlation function. Sub-Ohmic baths, which yield slow algebraic decay of coherence, correspond to smaller $\alpha$, while Ohmic or super-Ohmic baths, which produce faster decay, yield $\alpha$ closer to unity. Thus, $\alpha$ provides a compact and operational measure of bath-induced memory strength that connects the fractional master equation directly to physically observable decoherence profiles.
@@
Refer to caption Liouville $\alpha=1,\ \lambda=0$ Unitary, reversible Markovian Lindblad $\alpha=1,\ \lambda>0$ CPTP semigroup Non-Markovian QME $\alpha\!\neq\!1,\ \lambda>0$ Not always CPTP Fractional Lindblad $\prescript{C}{}{D}_{t}^{\alpha}\rho=\lambda\,\mathcal{L}\rho$ $0<\alpha<1,\ \lambda>0$ CPTP, power-law memory $\lambda\downarrow,~\alpha\uparrow$$\alpha\uparrowgeneral$ memory
Figure 1: Schematic hierarchy of quantum dynamical regimes unified by fractional subordination. The parameter $\alpha$ is the memory exponent of the renewal process governing the operational time of the dynamics: $\alpha=1$ yields Markovian Lindblad semigroups with exponential waiting times, while $0<\alpha<1$ generates long-tailed waiting statistics and non-Markovian evolution with algebraically decaying memory. The parameter $\lambda$ controls the dissipative strength of the underlying GKSL generator. Fractional or subordinated Lindblad dynamics ($0<\alpha<1$) therefore provide a CPTP bridge connecting unitary Liouville evolution ($\alpha=1,\lambda=0$), Markovian semigroups ($\alpha=1,\lambda>0$), and general memory–kernel quantum master equations ($\alpha\neq 1$).
@@
II.2 Fractional Master Equation
@@
II.2.1 Equivalence to convolution master equations
@@
The boundary term in Eq.,(11), $\frac{t^{\alpha-1}}{\Gamma(\alpha)}~\mathcal{L}\rho(0)$ arises generically when the Caputo form is converted into a differential convolution equation Chruscinski2017; Tarasov2021. For the fractional regime of primary interest in this work, $0<\alpha<1$, this term represents an initial-slip correction reflecting the mismatch between the initial state and the stationary state of the Lindbladian. Its behavior is fully analogous to short-time slip terms in NZ memory-kernel master equations and in CTRW models of anomalous relaxation Metzler2014. Although $t^{\alpha-1}$ diverges as $t\rightarrow 0$ for $0<\alpha<1$, its integrated influence is finite and it decays rapidly, contributing only transiently before the Mittag–Leffler relaxation dominates. By contrast, in the short-time Zeno regime analyzed later the effective order turns to be $\alpha=2$; in this case the same prefactor scales as $t^{\alpha-1}=t$ and therefore vanishes as $t\rightarrow 0$, so no divergence or anomalous enhancement occurs at early times.
@@
II.2.2 Mittag–Leffler solution and physical interpretation
@@ II.2.3 Fractional Subordination
The kernel satisfies $f_{\alpha}(u,t)\geq 0$ and $\int_{0}^{\infty}f_{\alpha}(u,t)~du=1$, which makes the evolution (17) represent a convex mixture of CPTP semigroups $e^{u\mathcal{L}}$ (detailed in Appendix C). Specifically, the subordination integral in Eq. (17) represents a probabilistic construction of fractional dynamics: the inverse–stable density $f_{\alpha}(u,t)\!\geq 0$ acts as a genuine probability distribution for the random operational time $u$ of an underlying stochastic clock. Consequently, $\Phi_{\alpha}(t)$ can be interpreted as the statistical mixture $\Phi_{\alpha}(t)=\mathbb{E}_{U(t)}[e^{U(t)\mathcal{L}}]$, i.e., an ensemble average of standard Markovian (Lindblad) evolutions sampled at random times $U(t)$ with power–law waiting statistics. This probabilistic representation ensures that complete positivity is preserved and explains how algebraic relaxation emerges naturally from the statistics of random operational time.
The fractional order $\alpha$ therefore acquires a clear operational meaning in the subordination picture: it determines the statistics of the random operational time $U(t)$. For an inverse–stable subordinator, the waiting-time distribution obeys $\psi_{\alpha}(\tau)\sim\tau^{-1-\alpha}$, so that $\alpha$ controls the heaviness of the tail and thus the persistence of memory in the stochastic time change. When $\alpha=1$, the process reduces to ordinary physical time with exponential waiting statistics, and one recovers the Markovian Lindblad semigroup $\{e^{t\mathcal{L}}\}$. When $0<\alpha<1$, the broad, power-law waiting-time statistics slow down the operational time and produce non-Markovian relaxation characterized by Mittag–Leffler or power-law decay. In this sense, $\alpha$ serves as a physically interpretable memory exponent rather than a phenomenological parameter.
[…]
Figure 2: General subordination replaces physical time by a random operational time. Fractional (inverse–stable) laws yield the fractional subordination, while other waiting-time distributions lead to non-fractional NZ kernels with short- or multi-scale memory (not necessarily power-law).
While inverse–stable subordination yields the fractional master equation, it is important to note that the subordination framework itself is far more general (see Figure 2). Any admissible waiting-time density $f(u,t)$ defines a subordinated evolution (17), and different choices of $f$ generate different classes of non-Markovian memory kernels Bochner1955; Feller1971; MeerschaertSikorskii2012. Fractional (inverse–stable) waiting-time laws are distinguished because they lead to a closed-form Caputo master equation, guarantee complete positivity for any GKSL generator, and produce Mittag–Leffler relaxation with algebraic memory Podlubny1999; Mainardi2000; BaeumerMeerschaert2001. However, subordination with non-stable waiting-time densities, such as exponential, stretched-exponential, or multi-exponential forms, remains mathematically valid and yields NZ-type memory kernels with short-ranged or multi-scale decay rather than power laws Breuer2002; Rivas2014; METZLER20001; Metzler_2004. Thus, fractional subordination should be viewed as a special, analytically tractable instance within a broader class of subordinated Lindblad evolutions: it targets systems whose non-Markovianity is dominated by algebraic or broad temporal correlations, while other waiting-time laws may be used to model qualitatively different memory structures.
The subordination representation (17) also has an important algorithmic interpretation. Because $e^{u\mathcal{L}}$ admits standard quantum–trajectory and quantum–jump unravelings Dalibard1992; Dum1992; Plenio1998, the subordination formula shows that any non-Markovian trajectory method can be extended to fractional, Markovian dynamics by replacing the physical time with a random clock $U(t)$ drawn from $f_{\alpha}(u,t)$, whose fluctuations become increasingly intermittent as $\alpha$ decreases. Smaller $\alpha$ thus leads to more strongly non-Markovian trajectory ensembles with longer sojourn periods between effective Lindblad updates. Thus, fractional dynamics corresponds to sampling ordinary Lindbladian trajectories at random operational times, with the memory effects arising entirely from the statistics of $U(t)$.
@@
Fractional Lindblad $\prescript{C}{}{D}_{t}^{\alpha}\rho=\mathcal{L}\rho$ Mittag–Leffler modes, long-memory NZ QME Laplace kernel: $\tilde{\kappa}_{\mathrm{NZ},\alpha}(s)=s^{1-\alpha}\mathcal{L}$ $(s-\tilde{\kappa}_{\mathrm{NZ},\alpha}(s))^{-1}$ HEOM Self-energy: $\Sigma(s)\sim s^{\chi}$ $(s^{\alpha}\mathbb{I}-\mathcal{L})^{-1}$,  $\alpha=\chi+1$ Subordination / CTRW $\rho(t)=\displaystyle\int_{0}^{\infty}f_{\alpha}(t,u)\,e^{u\mathcal{L}}\rho(0)\,du$ Lévy-time randomization Influence functionalQUAPI, MCTDH, path integralsexplicit bath correlations memory-kerneldeformationself-energydeformationoperational timerandomizationcoarse-grainingof long memory
Figure 3: Algebraic connections between the fractional Lindblad equation and mainstream non–Markovian approaches. Fractional dynamics arise as (i) a resolvent-level deformation of the NZ memory-kernel equation, (ii) a coarse-grained surrogate for HEOM self-energy structures, (iii) a subordinated Markovian process corresponding to Lévy-distributed operational times, and (iv) a compact representation of long-memory features also encoded in influence-functional methods such as QUAPI and MCTDH.
II.2.4 Algebraic connections to Nakajima–Zwanzig memory kernels and hierarchical equations of motion
The fractional master equation admits a unified algebraic structure that connects it directly to several mainstream non-Markovian formalisms. In particular, the Laplace-transform solution of the Caputo equation, Eq. (15), shows that fractional dynamics correspond to a spectral deformation of the standard Lindbladian resolvent $(s\mathbb{I}-\mathcal{L})^{-1}$ via the substitution $s\mapsto s^{\alpha}$. As discussed in the previous subsection, this deformation underlies the quantum analogue of subordinated Markov processes and continuous-time random walks (CTRWs) Bochner1955; Feller1971; BaeumerMeerschaert2001; Schilling2012Bernstein; MeerschaertSikorskii2012; Metzler_2004; METZLER20001, where heavy-tailed waiting-time statistics generate non-Markovian relaxation. We now elucidate how the same spectral deformation relates algebraically to other major approaches, including NZ projection-operator kernels, HEOM, and influence-functional methods.
Connection to NZ memory kernels.
A general NZ equation takes the convolution form Nakajima1958; Zwanzig1960
$\dot{\rho}(t)=\int_{0}^{t}\kappa_{\rm NZ}(t-\tau)\,\rho(\tau)\,d\tau,$ (23)
with Laplace-domain solution
$\tilde{\rho}(s)=(s\mathbb{I}-\tilde{\kappa}_{\rm NZ}(s))^{-1}\rho(0).$ (24)
Comparing this with the fractional resolvent (15) identifies the NZ kernel that reproduces fractional dynamics:
$\tilde{\kappa}_{{\rm NZ},\alpha}(s)=s^{1-\alpha}\mathcal{L}.$ (25)
In the time domain this corresponds to a weakly singular power-law kernel $\kappa_{{\rm NZ},\alpha}(t)=t^{\alpha-2}\mathcal{L}/\Gamma(\alpha-1)$, demonstrating that the fractional master equation is algebraically equivalent to a non-Markovian NZ equation with a completely monotone memory kernel.
Connection to HEOM.
In HEOM HEOM_1; HEOM_2; HEOM_3; HEOM_4, Gaussian environments generate a frequency-dependent self-energy $\Sigma(s)$ such that
$\tilde{\rho}(s)=(s\mathbb{I}-\mathcal{L}-\Sigma(s))^{-1}\rho(0).$ (26)
For structured spectral densities with low-frequency scaling $J(\omega)\sim\omega^{\chi}$, the bath correlation function acquires a power-law tail and the associated self-energy displays the asymptotic scaling $\Sigma(s)\sim s^{\chi}$. In this regime, the HEOM resolvent becomes algebraically equivalent to $(s^{\alpha}\mathbb{I}-\mathcal{L})^{-1}$ with $\alpha=\chi+1$, up to prefactors.
Connection to influence-functional approaches.
Path-integral methods such as QUAPI Makri1995QUAPI1; Makri1995QUAPI2, MCTDH-based open-system propagation Thoss2001MCTDH; Wang2003MCTDHReview, and related influence-functional techniques provide numerically controlled descriptions of non-Markovian dynamics by explicitly encoding bath correlation functions in discretized memory tensors or augmented Hilbert spaces. These approaches do not yield closed-form propagators and often scale exponentially with the bath memory time. The fractional master equation does not aim to reproduce their microscopic structure; instead, it serves as a compact surrogate for environments whose memory exhibits broad temporal support or algebraic tails, thereby complementing these methods at a coarse-grained level.
Together, these algebraic correspondences clarify the generality of fractional dynamics. For any GKSL generator, each Lindbladian eigenmode acquires a Mittag–Leffler relaxation profile, and the deformation $s\mapsto s^{\alpha}$ modifies the spectral response of the resolvent without altering the dissipator. Fractional dynamics therefore provide a low-parameter, CPTP surrogate that captures long-memory behavior reflected in NZ kernels, HEOM self-energies, and influence-functional approaches, while remaining algebraically consistent with established open-system theories.
III Numerical Demonstration and Discussion
To assess the physical relevance of the fractional Liouville–Lindblad dynamics, we benchmark it against a well-understood non-Markovian system for which an exact microscopic solution is available. The pure-dephasing spin–boson model provides an ideal testbed: its coherence dynamics are governed entirely by bath-induced memory effects, and the off-diagonal elements of the reduced density matrix admit a closed-form expression determined by the spectral density. This enables a precise, parameter-free comparison between (i) the exact coherence function derived from the full system–bath Hamiltonian (i.e. the ground-truth) and (ii) the coherence predicted by the fractional evolution obtained through subordination of a Lindblad generator, without numerical or model-dependent ambiguities. Our goal in this section is to evaluate whether the fractional propagation captures the characteristic non-exponential decay, long-memory behavior, and turnover between Gaussian and exponential regimes exhibited by the microscopic model. Establishing this agreement demonstrates that the fractional dynamics can faithfully reproduce structured non-Markovian behavior without requiring explicit microscopic modeling of the environment.
For this purpose, we now summarize the exact coherence dynamics of the pure-dephasing spin–boson model, which will serve as the microscopic reference against which the fractional evolution is compared. We work in the exactly solvable pure–dephasing limit of the spin–boson model,
@@ III Numerical Demonstration and Discussion
with independent boson modes $\{b_{k}\}$ in thermal equilibrium at inverse temperature $\beta$. Since $H,\sigma_{z}]=0$, populations are constants of motion; only the off–diagonal element (coherence) decays. If we assume a factorized state for the total spin-boson system
@@
as used in standard treatments of dissipative two-level systems and spin–boson dynamics Grifoni1998; Weiss2012. Table 1 provides the asymptotic scaling for the coherence that smoothly interpolates between Gaussian short-time decay, stretched-exponential-like intermediate behavior, and algebraic or saturating long-time tails, depending on $\chi$ and $t$ Leggett1987; Grifoni1998; Weiss2012; deVega2017. The real-time propagation of the dephasing functional $Q(t)$ and corresponding coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for $t\leq 0.2/\omega_{c}$ and $t\geq 5/\omega_{c}$ and three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are shown in Figure 4. In the weak-coupling and pure-dephasing limit, the reduced coherence $u(t)$ obeys a NZ convolution equation,
@@
coincides with the Gaussian-bath result in Eq. (29), confirming that $Q(t)$ fully characterizes the non-Markovian dephasing kernel. In the following sections, Eq. (35) serves as the baseline against which the conventional and fractional evolutions are evaluated.
@@
$Q(t)$ $\mathcal{O}\left(\frac{1}{2}\eta\,\Gamma(\chi+1)\,\omega_{c}^{2}\,t^{2}\right)$ $\mathcal{O}\left(C_{\chi}(\omega_{c})\,t^{1-\chi}\right)$ $\mathcal{O}\left(\eta\ln(\omega_{c}t)\right)$ $\mathcal{O}\left(Q_{\infty}-D_{\chi}\,t^{-(\chi-1)}\right)$
@@
III.1 Why Markovian dephasing misses $e^{-Q(t)}$?